Grade 12th passMechanics

the magnitude of two forces is such that when acting at right angles produces a resultant force of √20 and when acting at 60° produces a resultent of √28 . find out the nagnitude of two forces .

Varun srivastava
8 Years agoGrade 12th pass

3 Answers

Basma
8 Years ago
Let R be the resultant . R=root of A^2+B^2 +2ABcosx
20=A^2+B^2. When at angle 60,R=root of A^2+B^2+2ABcos60.  28=A^2+B^2+2AB/2
28=20+AB. AB=8
A^2+64/A^2=20
After calculations A=2,B=4
Devika
7 Years ago
The magnitude of two forces is such that when acting at right angle provide a resultant force of √20 &when acting at 60° produce a resultant =√28.Find the magnitude of the two forces
Atanu Maji
4 Years ago
The resultant of two forces p and q is at right angle of p so that the angle between the forces is cos inverse bracket minus p by q